OpenStudy (anonymous):

Which number is a solution of the inequality? x(7 – x) > 8 a) 2 b)8 c)-1 d)0

OpenStudy (anonymous):

i think a

OpenStudy (anonymous):

a

OpenStudy (anonymous1):

A, because 2(7 - 2) = 2*5 = 10.